What is the answer to -3b-2>1
Answer by jim_thompson5910(35256)

Step Statement Reason/Explanation
1. -3b-2 > 1 Start with the given inequality
2. -3b-2+2 > 1+2 Add 2 to both sides
3. -3b+0 > 3 Combine like terms (notice how the -2 and +2 add to 0)
4. -3b > 3 That zero goes away (since x+0 = 0+x = x)
5. -3b/(-3) < 3/(-3) Now divide both sides by -3 to isolate b. See Note below
6. b < -1 Divide


Note: when you multiply or divide both sides of an inequality by a negative number, the inequality sign will flip.

So the final answer is b < -1

In interval notation, the answer is